# Joint Estimation and Prediction of City-wide Delivery Demand: A Large Language Model Empowered Graph-based Learning Approach

### v1

@@ Abstract
The proliferation of e-commerce and urbanization has significantly intensified delivery operations in urban areas, boosting the volume and complexity of delivery demand. Data-driven predictive methods, especially those utilizing machine learning techniques, have emerged to handle these complexities in urban delivery demand management problems. One particularly pressing problem that has not yet been sufficiently studied is the joint estimation and prediction of city-wide delivery demand. To this end, we formulate this problem as a graph-based spatiotemporal learning task. First, a message-passing neural network model is formalized to capture the interaction between demand patterns of associated regions. Second, by exploiting recent advances in large language models, we extract general geospatial knowledge encodings from the unstructured locational data and integrate them into the demand predictor. Last, to encourage the cross-city transferability of the model, an inductive training scheme is developed in an end-to-end routine. Extensive empirical results on two real-world delivery datasets, including eight cities in China and the US, demonstrate that our model significantly outperforms state-of-the-art baselines in these challenging tasks.
1 Introduction
@@
Driven by the rapid expansion of e-commerce and the concomitant increase in urban freight traffic, urban delivery demand management is of paramount significance due to its profound implications for various aspects of urban life, including urban mobility, environmental sustainability, and economic efficiency, emerging as a critical area of inquiry in transportation research. The increasing volume of deliveries in metropolitan areas has exacerbated traffic congestion, leading to longer travel times, higher fuel consumption, and greater greenhouse gas emissions (Yannis et al., 2006). As e-commerce continues to proliferate, the volume and complexity of delivery operations in metropolitan areas have surged, necessitating more sophisticated and responsive management strategies. Thus, efficient management of delivery demand is crucial to alleviating these adverse effects, ensuring that delivery operations are both effective and sustainable (Black & Schreffler, 2010; Srinivas et al., 2022).
@@
To efficiently manage delivery demand in urban logistics systems, a variety of specific tasks are studied to improve operational efficiency and sustainability, such as parking time optimization (Reed et al., 2024), route planning (He et al., 2022), system architecture design (Azcuy et al., 2021), and strategic design (Janjevic et al., 2019). These tasks rely heavily on the accurate and timely estimation (Muñuzuri et al., 2012) and forecasting of the delivery demand (Hess et al., 2021; Nuzzolo & Comi, 2014). The significance of this issue is further underscored by several pressing challenges, including the high variability in demand patterns, the necessity for dynamic responsiveness in logistics operations, and precision in logistics planning. In this study, the term “delivery demand” is defined as the total number of pickup (or dropoff) delivery orders, including package and food orders, within a specified region of interest (ROI). Subsequently, the spatio-temporal evolution and distribution of city-wide demand patterns of all ROIs are examined.
@@
Despite great advances that have been made in many real-world problems (Zeng et al., 2019; Li et al., 2021; Wu & Wu, 2019; Gao et al., 2021; Hess et al., 2021; Liang et al., 2023), there are still some critical scenarios that remain underexplored and call for new modeling frameworks. Consider the case in which the demand values in operational depots are fully recorded in a region within a city. When new depots are planned for development in proximity to existing regions, managers endeavor to estimate and predict potential demands, as illustrated in Fig. 1. Furthermore, when a business seeks to expand to a new city where demand data are recorded over a relatively short period, it is essential to promptly evaluate the projected levels of future demand based on the limited records directly in order to ensure the adaptability of delivery strategies. In fact, operations in new regions have no historical data for reference. The model needs to exploit patterns from existing regions or transfer knowledge from the source cities to successfully predict expected demands.
@@
To address this challenging issue, we break down the complexity into three key aspects and will systematically address each in a progressive manner.
@@
1. How to model intricate interactions between the demand patterns of existing depots and newly developed ones. It is anticipated that the construction of new warehouses will influence the demand evolution of existing depots. Concurrently, the operational sphere of influence of preexisting warehouses will similarly exert a synergistic effect on the demand for new warehouses. This leads to an intricate interaction between the two items and is crucial for accurate demand estimation. Therefore, we propose to capture these interactive phenomena through graph neural networks (GNNs). Using a structured graph representation, both region-specific patterns and region-wise dependencies are captured in a unified architecture.
@@
2. How to integrate adequate information to estimate unmeasured demands. The delivery demand is influenced by many associated factors (Fancello et al., 2017; Fabusuyi et al., 2020). In particular, geospatial factors play a pivotal role in the emergence of last-mile delivery demand (Suguna et al., 2021). In this sense, collecting sufficient geolocational data becomes a prerequisite for an accurate estimation of city demand. However, collecting such data is nontrivial and requires a laborious process. Additionally, different cities can have different data availability and source types. The specific geospatial data collected in a city may not be applicable to another city. A generally accessible data source is an urgent necessity. Fortunately, utilizing recent advances in Large language models (LLMs), we can extract general geospatial knowledge from the unstructured locational data by exploiting the context reasoning ability of them. By integrating the transferable knowledge from LLMs into the GNN architecture, our model can serve as a generalized demand estimator.
@@
3. How to guarantee the transferability of the model in new regions and cites. Last, operators and managers often confront with the “cold-start” scenario in delivery operation. Models overfit to the training demand distribution may fail to transfer to new regions or cities without sufficient historical data for model re-training and fine-tuning. Therefore, the zero-shot transferability can facilitate the dynamic adjustment and adaptability of the strategic design. To achieve this, we first construct a functional graph that is generalizable to different cities relying on the encoding of LLMs. Then the backbone GNN predictor is trained in an inductive manner with a joint reconstruction and forecasting task, ensuring its ability to generalize to new regions and cities in an end-to-end routine. Based on this strategy, our model achieves three levels of transfer: (1) new regions without any demand records; (2) new cities with full current demand observations; and (3) new cities with partial current demands.
@@
We formalize the city-wide delivery demand joint estimation and prediction problem as a graph-based learning model and encourage transferability by an inductive training scheme;
We integrate the geospatial knowledge encoding from LLMs into the model, enhancing the capture of both region-specific demand patterns and region-wise interactions;
Comprehensive evaluations using real-world urban delivery order datasets containing 8 cities from both China and the US demonstrate the effectiveness of our model.
@@ 2 Related Work
2.1 Urban Delivery Demand Management
The significant increase in urbanization and the advent of online commerce have elevated delivery demand management to a central concern in the realm of urban logistics systems. To evaluate the operational performance of the existing warehouse and delivery strategy, the records of pickup and delivery orders are collected and processed to develop analytical and data-centric models. These models are used to address real-world problems associated with logistics demand management, including last-mile delivery route planning (Zeng et al., 2019; Li et al., 2021), time of arrival estimation (Wu & Wu, 2019; Gao et al., 2021), route prediction (Wen et al., 2022), and demand prediction (Hess et al., 2021; Liang et al., 2023; Nuzzolo & Comi, 2014). The majority of existing studies focus on either package delivery or food delivery processes, utilizing private data within a singular logistics system. This increases the probability of drawing a conclusion that is biased towards a specific dataset. It is fortunate that an open-source dataset from industry (Wu et al., 2023) considerably facilitates data-driven applications. In addition to the above-studied problems, another fundamental problem is to predict the demand level of newly developed depots. The issue is further complicated by the lack of historical data on the performance of new warehouses, as well as the impact of existing warehouses on the operations of new businesses.
@@ 2.2 Large Language Models for Spatial-Temporal Data
In recent years, large language models (LLMs) have made significant strides in the field of natural language processing. The advent of generative pre-trained transformers (GPTs) has ushered in a new era of modern machine learning, where large models have the potential to become a foundational architecture for a range of general language-related tasks. Representative LLMs include BERT (Devlin et al., 2018), Llama (Touvron et al., 2023), and Mistral (Jiang et al., 2023), and have been widely adopted for a wide range of disciplines. LLMs are distinguished for their excellent reasoning ability and in-context learning capability, which allows for few-shot or even zero-shot tasks (Brown et al., 2020; Kojima et al., 2022; Wei et al., 2021). A large number of studies have devoted to explore the potentials of LLMs to address fundamental problems in various areas. Most related to the problem studied, LLMs have demonstrated effectiveness in time series forecasting (Jin et al., 2023a; Garza & Mergenthaler-Canseco, 2023; Jia et al., 2024; Chang et al., 2023), spatial-temporal forecasting (Li et al., 2024; Jin et al., 2023b; Zhang et al., 2024), and geospatial knowledge mining (Gurnee & Tegmark, 2023; Manvi et al., 2023; 2024; Tang et al., 2024). Moreover, LLMs have shown the potential to be integrated into delivery modeling, such as route optimization (Liu et al., 2023; Qu et al., 2023). Most of these methods adopt LLMs as backbone predictors either by transforming the input domain data to language tokens or fine-tuning LLMs with domain-specific instructions. In contrast to the aforementioned approach, we utilize LLMs as a generic geospatial encoder and enhance the downstream specialized model through the incorporation of LLM embedding, as opposed to employing them as the fundamental architectural component.
@@ 2.3 Spatial-Temporal Graph Neural Networks
Spatial-temporal graph neural networks (STGNNs) have become the leading approaches for graph-based spatiotemporal forecasting. STGNNs adopt sequential techniques such as recurrent neural networks (RNNs), temporal convolutional networks (TCNs) and self-attention to process temporal correlations. Spatial techniques such as graph convolution, node embedding, and message-passing neural networks are employed for exploiting spatial relational interactions. STGNNs have shown promising results using them for traffic flow forecasting, air quality prediction, and energy consumption management (Yu et al., 2018; Li et al., 2018; Wu et al., 2019b; Bai et al., 2020; Guo et al., 2019; Zheng et al., 2020; Shao et al., 2022c; Cini et al., 2022b).
A recent trend has shifted the focus from the design of advanced architectures to the manipulation of data-specific patterns. Learnable node embedding is a representative solution (Shao et al., 2022a; Cini et al., 2023b; Nie et al., 2023a). Learnable embeddings are fed into modules of STGNNs and learned end-to-end with the forecasting tasks, enhancing the ability to identify node-specific patterns. Specifically, Shao et al. (2022b) used learnable embeddings for all nodes, time of day, and day of week stamps, allowing MLPs to achieve competitive performance in several datasets. Cini et al. (2023b) offered a systematic interpretation of node embedding as local effects and integrated it into a global-local architectural template. Fourier positional encoding was adopted in (Nie et al., 2024b) to enable canonical MLPs to capture high-frequency structures in traffic data.
Apart from the spatiotemporal forecasting tasks, STGNNs have also demonstrated effectiveness in other important tasks, such as the time series and spatiotemporal data imputation (Liang et al., 2022; Cini et al., 2022a; Marisca et al., 2022; Nie et al., 2024a), spatiotemporal kriging (Wu et al., 2021b; c; Nie et al., 2023c; Zheng et al., 2023), and graph-based virtue sensing (De Felice et al., 2024). As we aim to forecast demand patterns for both operating areas with historical records and new areas without histories, the joint demand estimation and forecasting problem studied in this paper lies at the intersection of the two research venues, needing an integrated solution.
@@ 3.1 Problem Conceptualization and Formulation
which denote the intra-location and inter-location covariate sets, respectively. denotes the correlation between observed regions, which can be used to learn the relations between observations to predict future values. This is frequently discussed in the time series forecasting setting. includes the latent interactions between unobserved variables, which can be inferred from the message propagation of latent variables. Finally, learning in is the key to estimating the unmeasured demands for new regions. In this case, a similarity metric is needed to describe the relations between observed and unobserved locations. For example, proximity between coordinates can be used as a proxy in kriging models (Wu et al., 2021b). Instead, we model these dependencies using the functional similarity based on the knowledge extracted from LLMs. Meanwhile, we also associate each location with an encoding derived from LLMs, serving as an additional covariate for .

### v3

@@ Abstract
The proliferation of e-commerce and urbanization has significantly intensified delivery operations in urban areas, boosting the volume and complexity of delivery demand. Data-driven predictive methods, especially those utilizing machine learning techniques, have emerged to handle these complexities in urban delivery demand management problems. One particularly pressing issue that has yet to be sufficiently addressed is the joint estimation and prediction of city-wide delivery demand, as well as the generalization of the model to new cities. To this end, we formulate this problem as a transferable graph-based spatiotemporal learning task. First, an individual-collective message-passing neural network model is formalized to capture the interaction between demand patterns of associated regions. Second, by exploiting recent advances in large language models (LLMs), we extract general geospatial knowledge encodings from the unstructured locational data using the embedding generated by LLMs. Last, to encourage the cross-city generalization of the model, we integrate the encoding into the demand predictor in a transferable way. Comprehensive empirical evaluation results on two real-world delivery datasets, including eight cities in China and the US, demonstrate that our model significantly outperforms state-of-the-art baselines in accuracy, efficiency, and transferability. PyTorch implementation is available at: https://github.com/tongnie/IMPEL.
keywords:
Urban logistics, Delivery demand, Demand estimation, Graph-based forecasting, Large language models1 Introduction
@@
Driven by the rapid expansion of e-commerce and the concomitant increase in urban freight traffic, urban delivery demand management is of paramount significance due to its profound implications for various aspects of urban life, including urban mobility, environmental sustainability, and economic efficiency, emerging as a critical area of inquiry in transportation research. The increasing volume of deliveries in metropolitan areas has exacerbated traffic congestion, leading to longer travel times, higher fuel consumption, and greater greenhouse gas emissions (Yannis et al., 2006). As e-commerce continues to proliferate, the volume and complexity of delivery operations in metropolitan areas have surged, necessitating more sophisticated and responsive management strategies. Thus, efficient management of delivery demand is crucial to alleviating these adverse effects, ensuring that delivery operations are both effective and sustainable (Black and Schreffler, 2010; Srinivas et al., 2022).
@@
To efficiently manage delivery demand in urban logistics systems, a variety of specific tasks are studied to improve operational efficiency and sustainability, such as parking time optimization (Reed et al., 2024), route planning (He et al., 2022), system architecture design (Azcuy et al., 2021), and strategic design (Janjevic et al., 2019). These tasks rely heavily on the accurate and timely estimation (Muñuzuri et al., 2012) and forecasting of the delivery demand (Hess et al., 2021; Nuzzolo and Comi, 2014). The significance of this issue is further underscored by several pressing challenges, including the high variability in demand patterns, the necessity for dynamic responsiveness in logistics operations, and precision in logistics planning. In this study, the term “delivery demand” is defined as the total number of pickup (or dropoff) delivery orders, including package and food orders, within a specified region of interest (ROI). Subsequently, the spatio-temporal evolution and distribution of city-wide demand patterns of all ROIs are examined.
@@
Despite great advances that have been made in many real-world problems (Zeng et al., 2019; Li et al., 2021; Wu and Wu, 2019; Gao et al., 2021; Hess et al., 2021; Liang et al., 2023a), there are still some critical scenarios that remain underexplored and call for new modeling frameworks. Consider the case in which the demand values in operational regions are fully recorded in a region within a city. When new regions are planned for development in proximity to existing regions, managers endeavor to estimate and predict potential demands, as illustrated in Fig. 1. Furthermore, when a business seeks to expand to a new city where demand data are recorded over a relatively short period, it is essential to promptly evaluate the projected levels of future demand based on the limited records directly in order to ensure the adaptability of delivery strategies. In fact, operations in new regions have no historical data for reference. The model needs to exploit patterns from existing regions or transfer knowledge from the source cities to successfully predict expected demands.
@@
To address this challenging issue, we break down the complexity into three key aspects as follows.
@@
1. How to model intricate interactions between the demand patterns of existing regions and newly developed ones. It is anticipated that the construction of warehouses in new regions will influence the demand evolution of existing regions. Concurrently, the operational sphere of influence of preexisting warehouses will similarly exert a synergistic effect on the demand for new warehouses. This leads to an intricate interaction between the two items and is crucial for accurate demand estimation and prediction. Therefore, we assume that these interactive phenomena are composed of both region-specific patterns and region-wise dependencies, and propose to capture them using individual-collective graph representations.
@@
2. How to integrate adequate information to estimate unmeasured demands. The delivery demand is influenced by many associated factors (Fancello et al., 2017; Fabusuyi et al., 2020). In particular, geospatial factors play an essential role in the emergence of last-mile delivery demand (Suguna et al., 2021). In this sense, collecting sufficient geolocational data becomes a prerequisite for an accurate estimation of city demand. However, collecting such data is nontrivial and requires a laborious process. In addition, different cities can have different data availability and source types. Specific geospatial data collected in a city may not be applicable to another city. An accessible data source is an urgent necessity. Fortunately, utilizing recent advances in large language models (LLMs), we can extract general geospatial knowledge from the unstructured locational data by exploiting the context reasoning ability of them. By integrating the transferable knowledge from LLMs into the spatial-temporal graph neural network (STGNN), our architecture can serve as a generalized demand estimator.
@@
3. How to guarantee the transferability of the model in new regions and cites. Last, operators and managers often confront with the “cold-start” scenario in delivery operation. Models overfit to the training demand distribution may fail to transfer to new regions or cities without sufficient historical data for model re-training and fine-tuning. Therefore, the zero-shot transferability can facilitate the dynamic adjustment and adaptability of the strategic design. To achieve this, we first construct a functional graph that is generalizable to different cities relying on the encoding of LLMs. Then the backbone STGNN predictor is trained in an inductive manner with a joint reconstruction and forecasting task, ensuring its ability to generalize to new regions and cities in an end-to-end routine. Since the LLM-generated encoding requires no training and can be obtained offline, it enables prediction with varying numbers of new regions and cities at different scales. Based on this strategy, our model achieves three levels of transfer: (1) new regions without any demand records; (2) new cities with full current demand observations; and (3) new cities with partial current demands.
@@
1.
We highlight the significance of location-based modeling in delivery demand joint estimation and prediction task and develop a data-driven scheme for extracting geolocation knowledge from LLMs;
2.
We present a method for integrating geolocation encoding into graph-based deep learning architectures. This approach allows for the modeling of both region-specific patterns and region-wide interactions in a unified graph representation, while simultaneously preserving the transferability across cities.
3.
Comprehensive evaluations using real-world urban delivery order datasets containing 8 cities from both China and the US demonstrate the effectiveness of our model. It is demonstrated that the performance gains in accuracy of the encoding are generic and it substantially improves cross-city transferability.
@@ 2 Related Work
2.1 Data-driven Urban Delivery Demand Management
The significant increase in urbanization and the advent of online commerce have elevated delivery demand management to a central concern in the realm of urban logistics systems. To evaluate the operational performance of the existing warehouse and delivery strategy, the records of pickup and delivery orders are collected and processed to develop analytical and data-centric models (Wen et al., 2024). These models are used to address real-world problems associated with logistics demand management, including last-mile delivery route planning (Zeng et al., 2019; Li et al., 2021), time of arrival estimation (Wu and Wu, 2019; Gao et al., 2021), route prediction (Wen et al., 2022), and demand prediction (Hess et al., 2021; Liang et al., 2023a; Nuzzolo and Comi, 2014). The majority of existing studies focus on package delivery or food delivery processes, utilizing private data within a singular logistics system. This increases the probability of drawing a conclusion that is biased towards a specific dataset. It is fortunate that an open-source dataset from industry (Wu et al., 2023) considerably facilitates data-driven applications. In addition to the above-studied problems, another fundamental task is to predict the demand level of newly developed regions. The issue is further complicated by the lack of historical data on the performance of new regions, as well as the impact of existing warehouses on the operations of new businesses.
@@ 2.2 Large Language Models for Spatial-Temporal Data
In recent years, large language models (LLMs) have made significant strides in the field of natural language processing. The advent of generative pre-trained transformers (GPTs) has ushered in a new era of modern machine learning, where large models have the potential to become a foundational architecture for a range of general language-related tasks. Representative LLMs include BERT (Devlin et al., 2018), Llama (Touvron et al., 2023), and Mistral (Jiang et al., 2023), and have been widely adopted for a wide range of disciplines. LLMs are distinguished for their excellent reasoning ability and in-context learning capability, which allows for few-shot or even zero-shot tasks (Brown et al., 2020; Kojima et al., 2022; Wei et al., 2021). A large number of studies have devoted to explore the potentials of LLMs to address fundamental problems in various areas. Most related to the problem studied, LLMs have demonstrated effectiveness in time series forecasting (Jin et al., 2023a; Garza and Mergenthaler-Canseco, 2023; Jia et al., 2024; Chang et al., 2023), spatial-temporal forecasting (Li et al., 2024; Jin et al., 2023b; Zhang et al., 2024b), and geospatial knowledge mining (Gurnee and Tegmark, 2023; Manvi et al., 2023, 2024; Tang et al., 2024). Moreover, LLMs have shown the potential to be integrated into transportation and mobility research, such as delivery route optimization (Liu et al., 2023; Qu et al., 2023), traffic data imputation (Zhang et al., 2024a), mobility demand prediction (Wang et al., 2023; Liang et al., 2024a; Gong et al., 2024), individual travel routine generation (Wang et al., 2024). Most of these methods adopt LLMs as backbone predictors either by transforming the input domain data to language tokens or fine-tuning LLMs with domain-specific instructions. In contrast to the aforementioned approach, we utilize LLMs as a generic geospatial encoder and enhance the downstream specialized model through the incorporation of LLM embedding, as opposed to employing them as the fundamental architectural component.
@@ 2.3 Spatial-Temporal Graph Neural Networks
Spatial-temporal graph neural networks (STGNNs) have become the leading approaches for graph-based spatiotemporal forecasting. STGNNs adopt sequential techniques such as recurrent neural networks (RNNs), temporal convolutional networks (TCNs) and self-attention to process temporal correlations. Spatial techniques such as graph convolution, node embedding, and message-passing neural networks are employed for exploiting spatial relational interactions. STGNNs have shown promising results using them for traffic flow forecasting, air quality prediction, and energy consumption management (Yu et al., 2018; Li et al., 2018; Wu et al., 2019a; Bai et al., 2020; Guo et al., 2019; Zheng et al., 2020; Shao et al., 2022c; Cini et al., 2022).
A recent trend has shifted the focus from the design of advanced architectures to the manipulation of data-specific patterns. Learnable node embedding is a representative solution (Shao et al., 2022a; Cini et al., 2023b; Nie et al., 2024c). Learnable embeddings are fed into modules of STGNNs and learned end-to-end with the forecasting tasks, enhancing the ability to identify node-specific patterns. Specifically, Shao et al. (2022b) used learnable embeddings for all nodes, time of day, and day of week stamps, allowing MLPs to achieve competitive performance in several datasets. Cini et al. (2023b) offered a systematic interpretation of node embedding as local effects and integrated it into a global-local architectural template. Fourier positional encoding was adopted in (Nie et al., 2024b) to enable canonical MLPs to capture high-frequency structures in traffic data.
Apart from the spatiotemporal forecasting tasks, STGNNs have also demonstrated effectiveness in other important tasks, such as the time series and spatiotemporal data imputation (Liang et al., 2022; Cini et al., 2021; Marisca et al., 2022; Nie et al., 2024a; Xue et al., 2024), spatiotemporal kriging (Wu et al., 2021a, c; Nie et al., 2024c; Zheng et al., 2023), and graph-based virtue sensing (De Felice et al., 2024). As we aim to forecast demand patterns for both operating areas with historical records and new areas without histories, the joint demand estimation and forecasting problem studied lies at the intersection of the two research venues, needing an integrated solution.
Similar to the delivery demand estimation task in urban logistics system, recent studies have elaborated advanced STGNNs to address motility demand estimation problems in both micro and macro transportation systems (Xue et al., 2025). Representative works include bike sharing demand prediction (Liang et al., 2023c), dynamic flow distribution prediction in dockless e-scooter sharing system (He and Shin, 2020), trip generation for bike sharing platform planning (Liang et al., 2023b, 2024b), and mobility flow generation (Yao et al., 2020; Zhang et al., 2023; Liang et al., 2024c). Despite promising performance, these methods rely on handcrafted geospatial features with an arduous data collection process, such as POIs and distance metrics, to construct underlying graphs used as input to the model. Particularly, the transfer between observed locations and unobserved locations has been achieved in previous work by aggregating the representation learned for observed locations using a prescribed similarity metric (Liang et al., 2024b; Wei et al., 2024). However, this is sensitive to the choice of computation rules and the availability of supplementary data, making the model less generalizable across cities.
@@ 3.1 Problem Conceptualization and Formulation
where the unobserved historical demand is estimated and future demands for all regions are predicted. The standard demand time series forecasting task exploits rich historical patterns, and the demand estimation task emphasizes the relations between spatial units. The problem studied lies in the junction of the two. In contrast to previous work on demand estimation, we also advocate for the transferability of the model between different regions and cities, which is of high practical significance and yet to be fully explored.
@@
which denote the intra-location and inter-location covariate sets, respectively. denotes the correlation between observed regions, which can be used to learn the relations between observations to predict future values. This is frequently discussed in the time series forecasting setting. includes the latent interactions between unobserved variables, which can be inferred from the message propagation of latent variables. Finally, learning in is the key to estimating unmeasured demands for new regions. In this case, a similarity metric is needed to describe relations between observed and unobserved locations. For example, proximity between coordinates can be used as a proxy in kriging models (Wu et al., 2021a). Instead, we model these dependencies using the functional similarity based on the knowledge extracted from LLMs. Meanwhile, we also associate each location with an encoding derived from LLMs, serving as an additional covariate for .